OpenStudy (anonymous):

completely factor the polynomial. 5x^3+15x^2+5x+15=

OpenStudy (anonymous):

You factor it by finding the roots.

OpenStudy (anonymous):

well i got it to 5(x^3+3x^2+x+3) but now i'm not sure what to do next

OpenStudy (tkhunny):

Now try all the POSSIBLE RATIONAL values. +/- 1 and +/- 3

OpenStudy (anonymous):

not sure what you mean by that

OpenStudy (tkhunny):

If f(x) = x^3+3x^2+x+3, And if f(3) = 0, Then 3 is a zero of f(x) and x-3 is a factor of f(x). If f(x) = x^3+3x^2+x+3, And if f(-1) = 0, Then -1 is a zero of f(x) and x+1 is a factor of f(x).

OpenStudy (anonymous):

so x-3 works so would the answer be 5(x+3)(x^2+1)?

OpenStudy (anonymous):

That is good, so long as it can't be factored further.

OpenStudy (anonymous):

Thank you for helping :)